OpenStudy (anonymous):

solve the equation log (base 3) 9^x=-2 the solution set is _________ type an integer or simplified fraction (use comma to separate answers as needed).

OpenStudy (anonymous):

hey!!! you seem have problems with log? need my help?

OpenStudy (anonymous):

yes please! just learn them two days ago

OpenStudy (anonymous):

\[\log_{3} 9^x =-2---> 3^-2 = 9^x\]are you ok with that?

OpenStudy (anonymous):

what is it turn to? 3-2=9^x?

OpenStudy (anonymous):

that the formula, log (base a) b = c means a^c =b

OpenStudy (anonymous):

your a is 3 (base 3) your c is -2 (the whole log =-2), your b is 9^x

OpenStudy (anonymous):

oh okay i just didn't see what it turns to

OpenStudy (anonymous):

now, right side is 9^x = 3^2x is it ok?

OpenStudy (anonymous):

left side 3^-2 = 3^2x ---> 2x =-2 --> x =-1. is it ok?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

nothing to do, right?

OpenStudy (anonymous):

thank you! looked like a hard problem can you help with another one?

OpenStudy (anonymous):

ok post at a new one.